OpenStudy (mia99):

please help me find second derivative of 2y^3-xy=4

Nnesha (nnesha):

\[\huge\rm 2y^3-x\cdot y=4\]

OpenStudy (mia99):

i have the first dy/dx=\[y/6y ^{2}-x\]

OpenStudy (mia99):

the second derivative is kinda hard

OpenStudy (solomonzelman):

Example ?

OpenStudy (mathmale):

\[\huge\rm 2y^3-x\cdot y=4\]

OpenStudy (mathmale):

Assuming that you want the first derivative with respect to x, differentiate each term of the above equation with respect to x. For 2y^3 you must apply the Chain Rule, since y is assumed to be a function of x. Next, your x*y is a product; you must use the Product Rule, followed by the Chain Rule. What ist he derivative of a constant (4)?

OpenStudy (mia99):

0

OpenStudy (mathmale):

There's more, a lot more, to the first derivative of the given equation\[\huge\rm 2y^3-x\cdot y=4\]

OpenStudy (mathmale):

How would you differentiate this?